# Neural Wave Functions for Superfluids

### v3

@@ Abstract
Understanding superfluidity remains a major goal of condensed matter physics. Here we tackle this challenge utilizing the recently developed Fermionic neural network (FermiNet) wave function Ansatz [1] for variational Monte Carlo calculations. We study the unitary Fermi gas, a system with strong, short-range, two-body interactions known to possess a superfluid ground state but difficult to describe quantitatively. We demonstrate key limitations of the FermiNet Ansatz in studying the unitary Fermi gas and propose a simple modification that outperforms the original FermiNet significantly, giving highly accurate results. We prove mathematically that the new Ansatz, which only differs from the original Ansatz by the method of antisymmetrization, is a strict generalization of the original FermiNet architecture, despite the use of fewer parameters. Our approach shares several advantages with the FermiNet: the use of a neural network removes the need for an underlying basis set; and the flexibility of the network yields extremely accurate results within a variational quantum Monte Carlo framework that provides access to unbiased estimates of arbitrary ground-state expectation values. We discuss how the method can be extended to study other superfluids.
@@ I Introduction
The unitary Fermi gas (UFG) is a paradigmatic example of a strongly interacting system of two-component fermions that possesses superfluid ground states and lies in the crossover region between a Bardeen-Cooper-Schrieffer (BCS) superconductor and a Bose-Einstein condensate [2, 3]. The effective range of the interaction is zero and the $s$-wave scattering length diverges (the “unitarity limit”), so the UFG has no intrinsic length scale. The only remaining length is the inverse of the Fermi wavevector $1/k_{F}$, on which all thermodynamic quantities depend. For example, regardless of the particle density, the ground-state energy per particle of a unitary Fermi gas can be written as
@@
Because of the universality of the UFG model, it can be used to describe many real physical systems at different scales, such as the neutron matter in the inner crust of neutron stars [5] or quantum criticality of s-wave atomic superfluids [6, 7]. The size of the pairs in the UFG is comparable to the inter-particle spacing, which is also a feature of many high-$T_{c}$ superconductors [8, 9, 10]. As a result, the UFG has been studied extensively [11]. Although the UFG is an idealized model, it can be accurately realized in the laboratory using ultracold atomic gases in which the interactions have been tuned by using an external magnetic field to drive the system across a Feshbach resonance [12].
@@
The UFG has been studied for decades, but it remains difficult to calculate its ground-state properties accurately using analytic methods. Mean-field treatments such as BCS theory [13] give good results for systems with weak interactions, but there is no guarantee of success in the strongly interacting regime. As a result, various quantum Monte Carlo (QMC) methods [14, 15] have been used to simulate the properties of the UFG to high accuracy at zero and finite temperature. Methods used include variational Monte Carlo (VMC), fixed-node diffusion Monte Carlo (FN-DMC), fixed-node Green function Monte Carlo, auxiliary field Monte Carlo and diagrammatic Monte Carlo [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. However, a full quantitive description remains an open and challenging problem.
@@
Recent advances in machine learning algorithms and the growing availability of inexpensive GPU-based computational resources have allowed neural-network-based approaches to permeate many areas of computational physics, including lattice [27, 28, 29, 30] and continuum [1, 31, 32, 33] QMC simulations. Here we employ a neural network Ansatz within a VMC approach to study the unitary Fermi gas. The Ansatz we use, the Fermionic Neural Network (FermiNet) [1], gives very accurate results for atoms and molecules [1, 34, 35, 36] and has recently been applied to periodic solids and the homogeneous electron gas (HEG) with comparable success [37]. In the case of the HEG, the variational optimization of the FermiNet Ansatz discovered the quantum phase transition between the Fermi liquid and Wigner crystal ground states without external guidance [38]. In contrast, previous approaches required different Ansätze to be used for the two different phases. The FermiNet has not previously been applied to fermonic superfluids such as the UFG.
@@
The paper is organized as follows. Section II of this paper describes the architecture of the FermiNet. We find that the original FermiNet Ansatz is insufficient to capture the two-particle correlations of superfluids. This is the first example in which the FermiNet has been seen to fail both quantitatively and qualitatively. To remedy the problem, we modify the FermiNet architecture based on the idea of the antisymmetric geminal power singlet wave function (AGPs) [39, 40, 41, 42, 30], which we discuss in detail in Section III. The implementation of the AGPs wave function using the FermiNet, as well as its relation to the original block-diagonal multi-determinant FermiNet, are discussed in Section IV. Our computational results are presented in Section V, followed by a summary and discussion in Section VI. The Appendix includes detailed explanations and derivations of important formulae, as well as implementation and training details.
@@ II FermiNet
In conventional SJ Ansätze, the antisymmetry of the $N$-electron wave function is represented using Slater determinants, which are antisymmetrized products of single-particle orbitals. For simulations of solids it is common to use one determinant only; for molecules, a linear combination of determinants is usually employed. In both cases, the presence of determinants guarantees that the wave function has the correct exchange antisymmetry. To improve the representation of electronic correlations, especially the correlations that chemists call “dynamic”, the determinants are multiplied by a totally symmetric non-negative function of the electron coordinates known as a Jastrow factor. This acts to decrease the value of the wave function as pairs of electrons approach each other, reducing the total Coulomb repulsion energy.
If the Hamiltonian is independent of spin and all of the single-particle orbitals are eigenfuntions of total $S_{z}$, one can assign spins to the electrons and every Slater determinant can be factored into a product of spin-up and spin-down Slater determinants [43, 14]. The wave function is no longer antisymmetric under the exchange of electrons of opposite spin, but expectation values of spin-independent operators are unaltered. Including a spin-assigned Jastrow factor expressed in the form $e^{J}$, a one-determinant SJ Ansatz becomes:
@@
where $\alpha$ and $\bar{\alpha}$ are the two possible spin components of an electron, $r_{ij}^{\beta\alpha}=|\mathbf{r}_{i}^{\beta}-\mathbf{r}_{j}^{\alpha}|$, and $\eta_{\parallel}(r)$ and $\eta_{\nparallel}(r)$ are parameterized functions of a single distance argument. The coordinate-transformed SJ Ansatz is called a Slater-Jastrow-backflow (SJB) wave function and the new coordinates are called quasiparticle coordinates. Note that the quasiparticle coordinate $\mathbf{x}_{j}^{\alpha}$ is invariant under the exchange of any two position vectors in $\{\mathbf{r}_{/j}^{\alpha}\}=\{\mathbf{r}_{1}^{\alpha},\dots,\mathbf{r}_{j-1}^{\alpha},\mathbf{r}_{j+1}^{\alpha},\dots,\mathbf{r}_{{N}^{\alpha}}^{\alpha}\}$ or in $\{\mathbf{r}^{\bar{\alpha}}\}=\{\mathbf{r}_{1}^{\bar{\alpha}},\dots,\mathbf{r}_{N^{\bar{\alpha}}}^{\bar{\alpha}}\}$[14].
@@
The FermiNet [1] takes the idea of permutation equivariant backflow much further, replacing the orbitals $\phi_{i}^{\alpha}(\mathbf{r}_{j}^{\alpha})$ entirely by neural networks. The orbitals represented by these networks differ from SJB orbitals because they are not functions of a single three-dimensional vector $\mathbf{x}_{j}^{\alpha}$ but depend in a very general way on $\mathbf{r}_{j}^{\alpha}$ and all of the elements of the sets $\{\mathbf{r}_{/j}^{\alpha}\}$ and $\{\mathbf{r}^{\bar{\alpha}}\}$. They are best written as $\phi_{i}^{\alpha}(\mathbf{r}_{j}^{\alpha};\{\mathbf{r}_{/j}^{\alpha}\};\{\mathbf{r}^{\bar{\alpha}}\})$. The exchange antisymmetry is maintained because $\phi_{i}^{\alpha}(\mathbf{r}_{j}^{\alpha};\{\mathbf{r}_{/j}^{\alpha}\};\{\mathbf{r}^{\bar{\alpha}}\})$ is totally symmetric on exchange of any pair of coordinates in $\{\mathbf{r}^{\alpha}_{/j}\}$ or $\{\mathbf{r}^{\bar{\alpha}}\}$. Furthermore, because they are represented as neural networks, the FermiNet orbitals need not be expanded in terms of an explicit basis set, widening the class of functions they can represent [44]. In order to build functions with the correct exchange symmetry properties, a carefully constructed neural network architecture is used, which is described below.
@@
where $\mathbf{V}^{l}$ and $\mathbf{W}^{l}$ are matrices, $\mathbf{b}^{l}$ and $\mathbf{c}^{l}$ are vectors, and all of them are optimizable. The outputs from the final layer $L$ of the one-electron streams are used to build the many-particle FermiNet orbitals:
@@
where $\mathbf{w}^{k\alpha}_{i}$ is an optimizable vector and $g^{k\alpha}_{i}$ an optimizable scalar. The $\chi^{k\alpha}_{i}(\mathbf{r}^{\alpha}_{j})$ factor is an envelope function to ensure that the wave function satisfies the relavent boundary conditions. For example, in a system which requires the wave function to tend to zero as $|\mathbf{r}^{\alpha}_{j}-\mathbf{R}_{m}|\rightarrow\infty$, exponential envelopes are used:
@@
As mentioned earlier, FermiNet orbitals are not functions of one electron position $\mathbf{r}^{\alpha}_{i}$ only, but also depend on the positions of all of the other electrons in the system in an appropriately permutation invariant way. No Jastrow factor is needed as the electron-electron correlations are included in the network. The full FermiNet wave function is thus a determinant of the FermiNet orbitals $\phi^{k\alpha}_{i}(\mathbf{r}^{\alpha}_{j};\{\mathbf{r}^{\alpha}_{/j}\};\{\mathbf{r}^{\bar{\alpha}}\})$. Multiple determinants may also be used, in which case the wave function is a weighted linear combination:
$\Psi_{\text{FermiNet}}\left(\mathbf{r}^{\uparrow}_{1},\dots,\mathbf{r}^{\uparrow}_{N^{\uparrow}},\mathbf{r}^{\downarrow}_{1},\dots,\mathbf{r}^{\downarrow}_{N^{\downarrow}}\right)$
$=\sum_{k}\omega_{k}\det[\phi^{k\uparrow}_{i}(\mathbf{r}^{\uparrow}_{j};\{\mathbf{r}^{\uparrow}_{/j}\};\{\mathbf{r}^{\downarrow}\})]\det[\phi^{k\downarrow}_{i}(\mathbf{r}^{\downarrow}_{j};\{\mathbf{r}^{\downarrow}_{/j}\};\{\mathbf{r}^{\uparrow}\})].$ (11)
In practice, the weights $\omega_{k}$ are absorbed into the orbitals, which are not normalized.
The VMC method is then applied with the FermiNet Ansatz and the parameters of the network are optimized using a second-order method known as the Kronecker-factored approximate curvature algorithm [45]. The aim is to minimize the expectation value of the Hamiltonian $\langle H\rangle$, which acts as our loss function. For a more detailed explanation of the FermiNet architecture, see Pfau et al.  [1] and the discussion of the improved JAX implementation [46] in Spencer et al. [36].
@@
The FermiNet has only been used to study systems of electrons interacting via Coulomb forces to date, but can easily be adapted to systems of other spin-$1/2$ particles simply by changing the Hamiltonian. Here we use the periodic FermiNet Ansatz to approximate the ground state of the UFG Hamiltonian in a cubic box subject to periodic boundary conditions. Since there are no atomic nuclei and the wave function has no electron-nuclear cusps, the inputs to the one-electron streams are simpler than shown in Eq. (4), containing only the particle coordinates 111 Although, in principle, the particle coordinates $\mathbf{r}^{\alpha}_{i}$ are not needed in a periodic system, we have observed that including them often improves convergence and allows the variational Ansatz to achieve a lower energy. Hence, we have included them in all of our calculations. : $\mathbf{h}^{0\alpha}_{i}=\mathbf{r}^{\alpha}_{i}$
As will be demonstrated below, the original FermiNet is sufficient to learn the superfluid ground state for small systems but fails for large systems. Hence, we propose a modification to the method of building orbitals. The motivation for this modification comes from earlier work using antisymmetrized products of two-particles orbitals known as antisymmetrized geminal power (AGP) wave functions [39, 49, 50, 40, 42, 16, 30]. We describe the antisymmetrized geminal power singlet (AGPs) wave function in the next section.

### v4

@@ Abstract
Understanding superfluidity remains a major goal of condensed matter physics. Here we tackle this challenge utilizing the recently developed Fermionic neural network (FermiNet) wave function Ansatz [1] for variational Monte Carlo calculations. We study the unitary Fermi gas, a system with strong, short-range, two-body interactions known to possess a superfluid ground state but difficult to describe quantitatively. We demonstrate key limitations of the FermiNet Ansatz in studying the unitary Fermi gas and propose a simple modification based on the idea of an antisymmetric geminal power singlet (AGPs) wave function. The new AGPs FermiNet outperforms the original FermiNet significantly in paired systems, giving results which are more accurate than fixed-node diffusion Monte Carlo and are consistent with experiment. We prove mathematically that the new Ansatz, which only differs from the original Ansatz by the method of antisymmetrization, is a strict generalization of the original FermiNet architecture, despite the use of fewer parameters. Our approach shares several advantages with the original FermiNet: the use of a neural network removes the need for an underlying basis set; and the flexibility of the network yields extremely accurate results within a variational quantum Monte Carlo framework that provides access to unbiased estimates of arbitrary ground-state expectation values. We discuss how the method can be extended to study other superfluids.
@@ I Introduction
The unitary Fermi gas (UFG) is a paradigmatic example of a strongly interacting system of two-component fermions that possesses a superfluid ground state and lies in the crossover region between a Bardeen-Cooper-Schrieffer (BCS) superconductor and a Bose-Einstein condensate [2, 3]. The effective range of the interaction is zero and the $s$-wave scattering length diverges (the “unitarity limit”), so the UFG has no intrinsic length scale. The only remaining length is the inverse of the Fermi wavevector $1/k_{F}$, on which all thermodynamic quantities depend. For example, regardless of the particle density, the ground-state energy per particle of a unitary Fermi gas can be written as
@@
Because of the universality of the UFG model, it can be used to describe many real physical systems at different scales, such as the neutron matter in the inner crust of a neutron star [5] or the quantum criticality of an s-wave atomic superfluid [6, 7]. The size of the pairs in the UFG is comparable to the inter-particle spacing, which is also a feature of many high-$T_{c}$ superconductors [8, 9, 10]. As a result, the UFG has been studied extensively [11]. Although the UFG is an idealized model, it can be accurately realized in the laboratory using ultracold atomic gases in which the interactions have been tuned by using an external magnetic field to drive the system across a Feshbach resonance [12].
@@
The UFG has been studied for decades, but it remains difficult to calculate its ground-state properties accurately using analytic methods. Mean-field treatments such as BCS theory [13] give good results for systems with weak interactions, but there is no guarantee of success in the strongly interacting regime. As a result, various quantum Monte Carlo (QMC) methods [14, 15] have been used to simulate the properties of the UFG to high accuracy at zero and finite temperature. Methods used include variational Monte Carlo (VMC), fixed-node diffusion Monte Carlo (FN-DMC), fixed-node Green function Monte Carlo, auxiliary field Monte Carlo and diagrammatic Monte Carlo [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. However, a full quantitative description remains an open and challenging problem.
@@
Recent advances in machine learning algorithms and the growing availability of inexpensive GPU-based computational resources have allowed neural-network-based approaches to permeate many areas of computational physics, including lattice [27, 28, 29, 30] and continuum [1, 31, 32, 33] QMC simulations. Here we employ a neural network Ansatz within a VMC approach to study the unitary Fermi gas. The Ansatz we use, the Fermionic Neural Network (FermiNet) [1], gives very accurate results for atoms and molecules [1, 34, 35, 36] and has recently been applied to periodic solids and the homogeneous electron gas (HEG) with comparable success [37]. In the case of the HEG, the variational optimization of the FermiNet Ansatz discovered the quantum phase transition between the Fermi liquid and Wigner crystal ground states without external guidance [38]. In contrast, previous approaches required different Ansätze to be used for the two different phases. The FermiNet has not previously been applied to fermionic superfluids such as the UFG.
@@
The paper is organized as follows. Section II describes the architecture of the FermiNet. We find that the original FermiNet Ansatz is insufficient to capture the two-particle correlations of superfluids. Although a FermiNet wave function with one determinant and a sufficiently large neural network is in principle able to represent any fermionic state [1], it is often advantageous to use a network of a fixed size and a small linear combination of FermiNet determinants. In the case of the unitary Fermi gas, however, we find that the number of block-diagonal determinants required to describe the ground state accurately scales exponentially with the system size. This is the first example in which the FermiNet has been seen to fail both quantitatively and qualitatively, and suggests that the FermiNet wave function may not be able to represent arbitrary fermionic wave functions in practice. To remedy the problem, we utilize the neural-network part of the FermiNet architecture to build a different type of wave function based on the idea of an antisymmetric geminal power singlet wave function (AGPs) [39, 40, 41, 42, 30], which we discuss in detail in Section III. This leads to substantial improvements, even though the neural-network part of the wave function remains unchanged. The implementation of the AGPs wave function using the FermiNet, as well as its relation to the original block-diagonal multi-determinant FermiNet, are discussed in Section IV. Our computational results are presented in Section V, followed by a summary and discussion in Section VI. The Appendix includes detailed explanations and derivations of important formulae, as well as implementation and training details.
@@ II FermiNet
In conventional SJ Ansätze, the antisymmetry of the $N$-electron wave function is represented using Slater determinants, which are antisymmetrized products of single-particle orbitals. For simulations of solids, it is common to use one determinant only; for molecules, a linear combination of determinants is usually employed. In both cases, the presence of determinants guarantees that the wave function has the correct exchange antisymmetry. To improve the representation of electronic correlations, especially the correlations that chemists call “dynamic”, the determinants are multiplied by a totally symmetric non-negative function of the electron coordinates known as a Jastrow factor. This acts to decrease the value of the wave function as pairs of electrons approach each other, reducing the total Coulomb repulsion energy.
If the Hamiltonian is independent of spin and all the single-particle orbitals are eigenfunctions of total $S_{z}$, one can assign spins to the electrons and every Slater determinant can be factored into a product of spin-up and spin-down Slater determinants [43, 14]. The wave function is no longer antisymmetric under the exchange of electrons of opposite spin, but expectation values of spin-independent operators are unaltered. Including a spin-assigned Jastrow factor expressed in the form $e^{J}$, a one-determinant SJ Ansatz becomes:
@@
where $\alpha$ and $\bar{\alpha}$ are the two possible spin components of an electron, $r_{ij}^{\beta\alpha}=|\mathbf{r}_{i}^{\beta}-\mathbf{r}_{j}^{\alpha}|$, and $\eta_{\parallel}(r)$ and $\eta_{\nparallel}(r)$ are parameterized functions of a single distance argument. The coordinate-transformed SJ Ansatz is called a Slater-Jastrow-backflow (SJB) wave function, and the new coordinates are called quasiparticle coordinates. Note that the quasiparticle coordinate $\mathbf{x}_{j}^{\alpha}$ is invariant under the exchange of any two position vectors in $\{\mathbf{r}_{/j}^{\alpha}\}=\{\mathbf{r}_{1}^{\alpha},\dots,\mathbf{r}_{j-1}^{\alpha},\mathbf{r}_{j+1}^{\alpha},\dots,\mathbf{r}_{{N}^{\alpha}}^{\alpha}\}$ or in $\{\mathbf{r}^{\bar{\alpha}}\}=\{\mathbf{r}_{1}^{\bar{\alpha}},\dots,\mathbf{r}_{N^{\bar{\alpha}}}^{\bar{\alpha}}\}$[14].
@@
The FermiNet [1] takes the idea of permutation equivariant backflow much further, replacing the orbitals $\phi_{i}^{\alpha}(\mathbf{r}_{j}^{\alpha})$ entirely by neural networks. The orbitals represented by these networks differ from SJB orbitals because they are not functions of a single three-dimensional vector $\mathbf{x}_{j}^{\alpha}$ but depend in a very general way on $\mathbf{r}_{j}^{\alpha}$ and all the elements of the sets $\{\mathbf{r}_{/j}^{\alpha}\}$ and $\{\mathbf{r}^{\bar{\alpha}}\}$. They are best written as $\phi_{i}^{\alpha}(\mathbf{r}_{j}^{\alpha};\{\mathbf{r}_{/j}^{\alpha}\};\{\mathbf{r}^{\bar{\alpha}}\})$. The exchange antisymmetry is maintained because $\phi_{i}^{\alpha}(\mathbf{r}_{j}^{\alpha};\{\mathbf{r}_{/j}^{\alpha}\};\{\mathbf{r}^{\bar{\alpha}}\})$ is totally symmetric on exchange of any pair of coordinates in $\{\mathbf{r}^{\alpha}_{/j}\}$ or $\{\mathbf{r}^{\bar{\alpha}}\}$. Furthermore, because they are represented as neural networks, the FermiNet orbitals need not be expanded in terms of an explicit basis set, widening the class of functions they can represent [44]. In order to build functions with the correct exchange symmetry properties, a carefully constructed neural network architecture is used, which is described below.
@@
where $\mathbf{V}^{l}$ and $\mathbf{W}^{l}$ are matrices, $\mathbf{b}^{l}$ and $\mathbf{c}^{l}$ are vectors, and all of them are optimizable. We denote the number of hidden units in each layer in the one-electron stream by $n_{l}$ such that $\mathbf{h}^{l\alpha}_{i}\in\mathbb{R}^{n_{l}},l\in\{0,1,2,\dots,L\}$ 111 Although $n_{l}$ can be different in different layers, in practice, we use an equal number of hidden units in each layer except for the input layer, where $n_{0}$ is fixed by the number of particles in the system. . The outputs from the final layer $L$ of the one-electron streams are used to build the many-particle FermiNet orbitals:
@@
where $\mathbf{w}^{k\alpha}_{i}$ is an optimizable vector and $g^{k\alpha}_{i}$ an optimizable scalar. The $\chi^{k\alpha}_{i}(\mathbf{r}^{\alpha}_{j})$ factor is an envelope function to ensure that the wave function satisfies the relevant boundary conditions. For example, in a system which requires the wave function to tend to zero as $|\mathbf{r}^{\alpha}_{j}-\mathbf{R}_{m}|\rightarrow\infty$, exponential envelopes are used:
@@
As mentioned earlier, FermiNet orbitals are not functions of one electron position $\mathbf{r}^{\alpha}_{i}$ only, but also depend on the positions of all the other electrons in the system in an appropriately permutation invariant way. No Jastrow factor is needed as the electron-electron correlations are included in the network. The full wave function is thus a block-diagonal determinant (BD) of the FermiNet orbitals $\phi^{k\alpha}_{i}(\mathbf{r}^{\alpha}_{j};\{\mathbf{r}^{\alpha}_{/j}\};\{\mathbf{r}^{\bar{\alpha}}\})$. Multiple determinants may also be used, in which case the wave function is a weighted linear combination
$\Psi^{D}_{\text{Slater FermiNet}}\left(\mathbf{r}^{\uparrow}_{1},\dots,\mathbf{r}^{\uparrow}_{N^{\uparrow}},\mathbf{r}^{\downarrow}_{1},\dots,\mathbf{r}^{\downarrow}_{N^{\downarrow}}\right)$
$=\sum^{D}_{k}\omega_{k}\det[\phi^{k\uparrow}_{i}(\mathbf{r}^{\uparrow}_{j};\{\mathbf{r}^{\uparrow}_{/j}\};\{\mathbf{r}^{\downarrow}\})]\det[\phi^{k\downarrow}_{i}(\mathbf{r}^{\downarrow}_{j};\{\mathbf{r}^{\downarrow}_{/j}\};\{\mathbf{r}^{\uparrow}\})],$ (11)
where the superscript $D$ specifies the number of determinants of FermiNet orbitals in the linear combination of determinants that makes up the full wave function, and the “Slater FermiNet” subscript serves to specify this specific wave function Ansatz and is discussed in more detail below. In practice, the weights $\omega_{k}$ are absorbed into the orbitals, which are not normalized.
The VMC method is then applied to the FermiNet Ansatz and the parameters of the network are optimized using a second-order method known as the Kronecker-factored approximate curvature algorithm [46]. The aim is to minimize the expectation value of the Hamiltonian $\langle H\rangle$, which acts as our loss function. For a more detailed explanation of the FermiNet architecture, see Pfau et al.  [1] and the discussion of the improved JAX implementation [47] in Spencer et al. [36].
@@
The FermiNet has only been used to study systems of electrons interacting via Coulomb forces to date, but can easily be adapted to systems of other spin-$1/2$ particles simply by changing the Hamiltonian. Here we use the periodic FermiNet Ansatz to approximate the ground state of the UFG Hamiltonian in a cubic box subject to periodic boundary conditions. Since there are no atomic nuclei and the wave function has no electron-nuclear cusps, the inputs to the one-electron streams are simpler than shown in Eq. (4), containing only the particle coordinates 222 Although, in principle, the particle coordinates $\mathbf{r}^{\alpha}_{i}$ are not needed in a periodic system, we have observed that including them often improves convergence and allows the variational Ansatz to achieve a lower energy. Hence, we have included them in all of our calculations. : $\mathbf{h}^{0\alpha}_{i}=\mathbf{r}^{\alpha}_{i}$, with respect to an origin placed at one corner of the simulation cell. A detailed discussion of translational symmetry of the wave function can be found in section E of the Appendix.
As will be demonstrated below, the Slater FermiNet is sufficient to learn the superfluid ground state for small systems but fails for large systems. Hence, we propose a modification to the method of building orbitals. The motivation for this modification comes from earlier work using antisymmetrized products of two-particles orbitals known as antisymmetrized geminal power (AGP) wave functions [39, 50, 51, 40, 42, 16, 30]. We describe the antisymmetrized geminal power singlet (AGPs) wave function in the next section.
The authors of Ref. [1] used the term “FermiNet wave function” to refer to all wave functions constructed using a FermiNet neural network. Now that we are going to use almost the same neural network to generate AGPs-based pairing wave functions in addition, more precise terminology is required. Wave functions of the type introduced in Ref. [1], which contain many-particle generalizations of the one-particle orbitals that appear in Slater determinants, will be called one-determinant or multi-determinant Slater FermiNets. Wave functions built using determinants of many-particle generalizations of pairing functions will be called one-determinant or multi-determinant AGPs FermiNets. Since every AGPs FermiNet determinant is built using one pairing function or “geminal”, we also refer to one-geminal or multi-geminal AGPs FermiNets.